Adult tickets for a play cost $10 and child tickets cost $4. If there were 30 people at a performance and the theater collected $234 from ticket sales, how many adults and how many children attended the play?
Answer by checkley71(8403)

c+a=30 or a=30-c
10a+4c=234
10(30-c)+4c=234
300-10c+4x=234
-6c=234-300
-6c=-66
c=-66/-6
c=11 children tickets were sold.
30-11=19 adult tickets were sold.
proof
10*19+4*11=234
190+44=234
234=234